please help me solve these inequalities with solutions in interval notation:
[x+5]>=10
[x/2]>=8
[x]+5<=10
[x-5]<10
[2x]>=8
3[x-1]+7>25

Answer by edjones(8007)

|x+5|>=10 I assume we are talking about absolutes and inequalities.
.
x+5>=10
x>=5
.
x+5<=-10
x<=-15
.
x(-infinity,-15]U[5,infinity)
.
|x|+5<=10
|x|<=5
x<=5
x>=-5
x[-5,5]
.
You can do the rest.
.
Ed
|x-5|<10
x-5<10
x<15
.
x-5>-10
x>-5
x(-5,15)
.
Ed